Ahead: Week 8 of Increment 17.
ISS crew work cycle shifted another 30 min. to the left: wake-up 5:32am EDT; sleep 8:32pm (Shuttle crew 30 min later: 9:02pm).
Crew activities aboard the ISS centered on three major areas: (1) Spacewalk #3 (EVA-3), (2) more JLP (JEM Logistics Pressurized Module) outfitting, and (3) sample collections from Kibo air & surfaces plus Node-2 ITCS coolant.
Mission 1J's EVA-3 was completed successfully by Mike Fossum & Ron Garan in 6h 33min, accomplishing all its objectives.
(During the spacewalk, Fossum (EV1) & Garan (EV2) -
Removed the old NTA (Nitrogen Tank Assembly) at S1 truss,
Retrieved the new NTA from ESP-3 (External Stowage Platform 3),
Installed new NTA on S1, stowed the old NTA & connected NTA FQD (Fluid Quick Disconnect) jumper,
Collected two debris samples (not grease) at Port SARJ, using Kapton tape,
Re-installed the ETVCG (External TV Camera Group) camera, reworked with new power supply, at CP9 (Camera Port 9) on the left truss,
Removed thermal covers & launch locks from JEMRMS WVE/EVE (Wrist & Elbow Vision Equipment) cameras,
Removed JPM aft window launch locks,
Deployed fwd & aft MMOD (Micrometeoroid/Orbital Debris) shields on ACBM (Active Common Berthing Mechanism), and
Inspected & tightened a loose bolt on the fwd JTVE camera boom's JPM mounting which 'wobbled' yesterday while panning & tilting (late-added get-ahead task).
Official start time of the spacewalk was 9:55am EDT, about 37 minutes ahead of timeline, and it ended at 4:28pm. Total EVA duration (PET = Phase Elapsed Time) was 6h 33min.)
After wakeup at ~5:32am, ending the 8.5-hr sleep period before the spacewalk, the A/L CL (Airlock Crewlock) hatch was cracked (~6:57am) for a hygiene break/with mask prebreathe for Fossum & Garan, after spending the night on 10.2 psi campout. Around 7:22am, the hatch was closed again by IVs (Intravehicular Crewmembers) for EVA preparations in 10.2 psi, followed by EMU (Extravehicular Mobility Unit) purge & prebreathe. Sergey Volkov, Mark Kelly & Greg Chamitoff assisted the spacewalkers during EMU purge, EMU prebreathe and CL depressurization. EV1/EV2 egress followed after the EMUs were switched to batteries at 9:55am.
Prior to EVA start, CDR Volkov checked two Kodak DCS760 digital cameras with 28mm lens for the EVA (one with flash for the SARJ task) and verified deactivation of all onboard ham radio equipment (Kenwood in SM/Service Module, Ericsson in FGB) to prevent RF interference with the EMUs, as well as proper closure of the protective Lab window shutters.
Volkov serviced the Russian BMP (Harmful Impurities Removal System), starting the "bake-out" cycle to vacuum on absorbent bed #2 of the regenerable dual-channel filtration system. The regen process will be terminated tonight at ~7:50pm EDT. Filter bed #1 was regenerated yesterday. (Regeneration of each of the two cartridges takes about 12 hours and is conducted only during crew awake periods. The BMP's regeneration cycle is regularly done every 20 days.)
As part of regular preventive maintenance of RS (Russian Segment) ventilation systems, Sergey replaced the four dust filters (PF1-4) in the SM.
Preventive maintenance was also performed by FE-1 Oleg Kononenko in the Soyuz TMA-12/16S at the DC1 nadir port, cleaning the screen of the BVN fan/heater assembly in the Orbital Module (BO).
FE-2 Reisman worked on the ITCS (Internal Thermal Control System) in Node-2, using the new 1J-delivered CQMK (Coolant Quality Monitoring Kit) to collect return-to-ground fluid samples from the LTL (Low Temperature Loop) and MTL (Moderate Temperature Loop) sample ports after flushing (ammonia or OPA/Ortho-Phthalaldehyde test samples not required).
Later, Garrett floated into the Kibo laboratory and used the MAS (Microbial Air Sampler) kit to obtain microbiology air samples from mid-module, and the SSK (Surface Sample Kit) to collect/incubate microbiology samples from two prime surface sites for return to ground.
Sergey unstowed the two Russian pilot sighting instruments VP-2 & 'Puma' and installed them at SM window #8 for functional testing. (The Puma Portable Zoom Viewfinder is used to view remote objects and determine their angular position in the SM coordinate system in order to provide geographical reference of observed terrestrial objects, and to determine the target vector in a specified coordinate system. The 240K Pilot Sight (VP-2) is a collimator-type device for determining the direction to observed reference points relative to the station coordinate for geographical reference of observed terrestrial objects and to determine the direction vector to controlled & uncontrolled objects and measure their angular sizes.)
Oleg Kononenko conducted the periodic (currently daily) checkout/verification of IP-1 airflow sensors in the various RS (Russian Segment) hatchways, including the passageways PrK (SM Transfer Compartment) -ATV, PrK -RO (SM Working Compartment), PkhO (SM Transfer Tunnel) -RO, PkhO -DC1, PkhO -FGB PGO, FGB PGO -FGB GA, FGB GA -Node-1. (This checkup is especially important when the ventilation/circulation system has to cope with a larger crew on board, currently ten persons, and one of the two Russian SKV air conditioners off (SKV-1).)
Kononenko also continued the current work with the Russian KPT-2 science payload BAR-RM, Kelvin, Ira and TTM from the RSE1 laptop, started by Oleg last Friday. (Measurements are being taken in the RS, behind panels, near welds along SM structural rings and near the shell ring in the FGB for subsequent downlinking via BSR-TM channel. At the FGB pressurized shell ring, the cosmonauts were to inspect structural elements, equipment, and cable bundles for moisture, mold, or evidence of corrosion. The data, collected at locations such as feedthroughs, windows, hatches etc., are being used to get proficient in experimenting with ISS leak detection based on environmental data anomalies (temperature, humidity, and ultrasound emissions) at potential leak locations. The BAR-RM payload uses a remote infrared thermometer (Kelvin-Video), a thermohygrometer (Iva-6A), a heat-loss anemometer/thermometer (TTM-2), an ultrasound analyzer (AU-01), and a leak detector (UT2-03) to determine physical background signs of loss of ISS pressure integrity which could be indicative of leaks in the working compartments of the station. Measurements are taken in specific zones (13 in SM PkhO and 4 in DC1), both with lights, fans & ASU pump turned on and off.)
For hardware familiarization/handover, Sergey Volkov & Greg Chamitoff inspected and checked out the HMS RSP (Health Maintenance System/Respiratory Support Pack).
FE-2 Reisman performed calibration on the new CSA-O2 (CSA -Oxygen sensor) units #1043 & #1059 delivered on 1J. The old CSA-O2 instruments (#1041 & #1052) were packed for return to Earth.
Garrett, slated for return on the Discovery on 6/14, and MS5 Greg Chamitoff, who replaces him as ISS FE-2, had about 3.5 hrs scheduled for standard joint handover activities, to be continued through the docked period ahead.
MS1 Karen Nyberg & MS4 Aki Hoshide first operated the SSRMS (Space Station Remote Manipulator System) to 'fly' & support Ron Garan during the S1 NTA remove/install, then moved to the JLP container module to conduct Part 1B outfitting by removing CBM (Common Berthing Mechanism) hardware, e.g. CPAs (Controller Panel Assemblies), and installing avionics jumpers.
The FE-1 conducted the routine daily servicing of the SOZh system (Environment Control & Life Support System, ECLSS) in the SM. (Regular daily SOZh maintenance consists, among else, of checking the ASU toilet facilities, replacement of the KTO & KBO solid waste containers and replacement of EDV-SV waste water and EDV-U urine containers.)
The CDR completed the daily IMS (Inventory Management System) maintenance, updating/editing its standard 'delta file' including stowage locations, for the regular weekly automated export/import to its three databases on the ground (Houston, Moscow, Baikonur).
Aki & Karen spent several hours on cargo transfers from & to the Shuttle middeck.
The ISS crew completed their regular 2.5-hr. physical workout program (about half of which is used for setup & post-exercise personal hygiene) on the CEVIS cycle ergometer (CDR, FE-1, FE-2, FE-2-17), and TVIS treadmill (CDR, FE-1), RED resistive exercise device (FE-2). Day 2 of CEVS physical exercise for Greg Chamitoff.
Afterwards, Sergey copied the exercise data file to the MEC (Medical Equipment Computer) laptop for downlink, including the daily wristband HRM (Heart Rate Monitor) data of the workouts on RED, followed by their erasure on the HRM storage medium (done six times a week).
Preview of tomorrow's main activities: SM & ATV thruster firings for IWIS, crew conference & photo, JEM RMS final deployment, installation of two new BMCs (Battery Charger Modules) in rotated Airlock Avionics Rack (A/L1F1), JLP outfitting part 2 & ingress.
No CEO (Crew Earth Observation) photo targets uplinked for today.